OpenStudy (anonymous):

Find the curvature κ(t) of the curve X (t) = ( −4 sint )I + ( −4 sint ) j + ( 3 cost ) K ???

OpenStudy (anonymous):

make ur question clear

OpenStudy (anonymous):

This question is crystal clear k(t) is nothing more than 1/|v|(|dT/dt)| ill do a very easy example r(t)=(cos(t)i+(sin(t))j r'(t)=(-sin(t))i+(cos(t))j |r'(t)|=1 r'(t)/|r'(t)|=(-sin(t))i+(cos(t))j=T d/dt(T)=(-cos(t))i+(-sin(t))j |dT/dt|=1 k(t)=1/1(1)=1